OpenStudy (anonymous):

Two trains leave the station at the same time. The first train is traveling north and the second train is traveling south 20 mph slower. After 2 hours, they are 180 miles apart. How fast is each train moving?

OpenStudy (anonymous):

Please also tell me the steps to figuring this out as well :)

OpenStudy (anonymous):

ok so i think i got rthis

OpenStudy (anonymous):

43 If a-b=7 and a and b are both positive integers, what is the minimum possible value of a+b [0, 9, 8, 11] 45. Tiffany, who is president of the senior class at her high school, sold $67.50 worth of tickets for the music festival. If she sold 20 more student tickets than adult tickets, and the cost for student tickets was 75 cents and adult tickets $1.00, how many of each kind of ticket did he sell? 46. A high school play grossed $122.50. If there were twice as many student tickets sold as there were adult tickets, and if each student ticket sold for 50 cents and each adult ticket for 75 cents, how many students and how many adult tickets were sold for the play? A47. Eighty-six coins made up of dimes and quarters amount to $14.60. How many of each of these coins are there? 48. A piggy bank contained 110 coins including nickels, dimes, and quarters. If there are 20 more nickels than quarters, and the number of dimes is equal to the number of quarter, how many of each kind of coin are there if the total collection is $13.00? 49. A news boy collected $68.50 on his paper route one week. The collection consisted of nickels, dimes, and quarters. If he had 70 more nickels than quarters and 20 more dimes than nickels. How many of each kind of coin did he have? 50. Angela has $11.25 in nickels, dimes, and quarters. She has 3 times as many nickels as dimes and 5 more quarters than dimes. How many of each kind of coin does she have? 51. A student a Ryan High has a collection of quarters and nickels amounting to $18.70. If there are 20 more nickels than quarters, how many of each kind of coin does the student have? 52. A man has $10.50 in change consisting of twice as many half dollars as quarters and five times as many dimes as quarters. How many half dollars, quarters, and dimes does she have? 53. In a collection of coins that has a value of $4.90, there are 20 more nickels than dimes. How many nickels and dimes are there in the collection? 54. Ratios: write the following ratios and reduce to lowest terms: (330 to 99, 16 ft to 4 yds, 300 m to 1 km, 12.5 gal to 2 qt) 55. Golf balls formerly sold at $1.00 each. Now on sale, they sell for $9.00 a dozen. What is the ratio of the former price to the sale price? 56. A jar contains 5 white marbles, 6 black marbles and 4 green marbles. If Alan removes 3 white marbles, 2 black marbles, and 1 green marble, what is the probability that the next marble removed at random is black? [4/5, 4/9, 1/3, 6/15, 4/15] 57. What is the ratio of the perimeter of square A to the square B? 58. What is the ratio of the perimeter of rectangle C to the rectangle D? 59. What it the ratio of the perimeter of triangle E to the triangle F? ______ _________ _____ _______ 14 cm Square 6 rectangle 9 8 cm triangle A57 57 A58 A58 6 10 18 cm _10 in_ __10 ft A59 A59 ____18 in__ _____20 ft___ ____12 cm___ Fabric Cost A 3 yards for $8 4x A61 B 2 yards for $6 3x (4x,3x) A60 C 4 yards for $9 2x D 5 yards for $7 x E 6 yards for $5 0 1x 2x 3x 4x 5x X A60. Which fabric costs the least per square yard? A61. If the area of triangle is 54, then x=? 62. What is the width of a rectangle with area 48 x2 and a length of 24x? 63. 103 (105 +105 ) 104 64. If x=b+4 and y=b-3, then in terms of x and y, b=? 65. If x= 1 / y +1 and y≠1, the y=? 66. If ∆(x)=x+1 and ⌂(x)=x-1, then which of the following is equal to ∆(3) x ⌂(5)? [⌂(17), ⌂(20), ∆(8), ∆(14)] y (a,a) (b,a) A67. What is the area of the shaded portion? (o,a) ___________________ Expressed in terms of a and b? [a(a-b), a(b-a), b(a-b), b(b-a), ab] A67 __________ (a,0)_ ___ (b,0 ) _x 68. +1, 0, -1, +1, 0, -1, +1,0,-1,... and so on, where the last number is +1, has a sum of [0, -2, -1, +1, +2] 69. Jesse standing at point A walks 1 mile south, 3 miles east, 5 miles north, and 7 miles west to reach pint B. How many miles of the length of line segment AB ? 70. Express as a ratio the number of 22-cent stamps you can buy if you have C cents. 71. Mr. Ho gets $6 an hour for doing 8 hours work, and 1.5 times that for overtime. How much does he earn if he works 12 hours? 72. Find x:│4x-2│=6, │3-0.5y│=-7, │-0.75m│=9, 2│x+7│=12 73. Four out of five students at West High take a mathematics course. If the enrollment at West is 785, how many students take mathematics? 74. (ratios and Proportions) find the appropriate solutions; (4:n=8:5, n:2/3=n/72, n:12=3:4 -n:6=-6:12) 75. Ratios: write the following ratios and reduce to lowest terms: (330 to 99, 16 ft to 4 yds, 300 m to 1 km, 12.5 gal to 2 qt) _________ ________________________ ________ 1960 ________________________ A76 ________________________ 20" A77 1970 ___________________A78 __10cm__ ________ ________________________ 14" 1980 _______________________ 1 2 3 4 5 6 7 8 A76. What is the area of the shaded triangle in this square (one side 10 cm)? A77. An arched window is in the shape of a rectangle with a semicircle on top. If the base of the rectangle is 14 inches and the height of the rectangle is 20 inches ,find the perimeter of the whole window ? [280 in, 68 in, 76 in, 77 in, 56 in] A78. The bar graph (population in thousands from 1 to 8) shows the population of a town during a 30-year period. By what number of people, did the population decrease between 1960 and 1970" ? [2,500 1,500 15,000 1,000] 79. How much greater is the volume of a cube whose edge is 5 inches than the volume of a cube whose edge is 4 inches? A80. There are 9 tiles with single-digit numbers painted on them. If xy is the product of the value of any tow tiles selected at random, how many different possible values for xy are there? □ □ □ □ □ □ □ □ □ .A80 [4,7,10, 12, 16] 0 1 0 2 0 3 0 40 . 81. (5/16) x ( ½ ) +3/8 =? 82. Using exponents, write 359 in expanded form. 83. At a factory, three out of every 1,000 parts produced are defective. In a day, the factory can produce 25,000 parts. How many of these parts would be defective? 84. (Inequalities) 3m+2<7 [m≤5/3, m≤2, m<2, m>2, m<5/3] 85. -6<2S/3+6≤2 [-6<S≤-3, S≥-6. -18<S≤-6, -2<S≤-4/3, S≤-6] 86. A pack of 6 potatoes weighed 3.75lbs. What was the average weight of each potato? 87. Write as percent: [9 hundredths, ½ hundredth, 600 hundredths, 260/100] 88. Solve the following polynomials: 3 x3y2-4xz -6x 3y2 =? 11q 2 r +4q2r +8q2 r =? 5p2t x 3 p2t=? (2a+b)(3a2 +ab+b2 )=? . 89. (x2+x-6) divided by (x-2)=? (m2+m-14) divided by (m+4)=? 90. (Simplifying)16b2-25z2=? X2-2x-8=? m2p2 +mpq +6q2=? 2h3+2h2t+4ht2 A91. Operations with Polynomials: 9a2 b+3c+2a2 b+5c=? 3x+2y+16x+3z+6y=? 3ac2+2b2c+7ac2+ 2ac2+b2 c=? 92. 5p2t x 3p2t=? (4m + p)(3m-2p)=? (2a+b)(3a2+ab+b2)=? (6t2+2t+1) 3t=? 93. The posts of a fence are 6 feet apart. The distance from the first post to the last post is 126 feet. How many posts are in the fence? 94. In an electrical device, the gap between 2 metal prongs is .002 inch when the temperature is 30°. If each of the metal prongs increases in length at a rate of 1/10,000 inch for each 1°F increase in temperature, at what temperature will the prongs make contact? 95. If a grams of substance 1 is mixed with b grams of substance 2, the percent of substance 1 in the mixture is [100a/b, 100b/ a+b, 100a/ a+b, a+b / 100a] 96. If p#q = 2 (p+2 q), then if a #b = b # a, which of the following must be true?[ a=b, a=-b, a=b=c, a=2b] 97. If 12 pencils cost 42 cents, how much would 100 pencils cost? 98. It takes a crew of 6 man to wash 84 windows of a hospital in one day. At the same rate how many men would it take to wash 140 windows? 99. If it takes 10.5 bushels of grapes to make 30 quarts of wine, how many bushels of grapes will lit take to make 3,200 quart of wine? 100. If 6 flowerpots cost $17.94, how much would 20 flowerpots cost? * Quantitative comparisons* : Mark A if the quantity in Column A is greater Mark B if the quantity in Column B is greater Mark C if the two quantity are equal Mark D if the relationship can't be determined from the information given Note: Information that applies to both quantities is centered above both columns A symbol that appears in both columns means the same thing in both columns ___________Column A _____________Column B________________ A. (1+2)-(3+4) ( 1/5 + 25) - ( 3/5+4/5) 5 2 Given 2x = y where x>z and y>0 B. x - y y - x Given the numbers 6,8,12,13,16 C. The mean of these numbers The median of these numbers B1. What is -27 subtracted from -36, divided by nine? B2. Simplify [8+(-4) ]x (-5-7)=? B3. What are the expresses of the number in scientific notation ? ( 185,000 and 0.00352) B4. (4 x 103) + (8 x 102) + (6 x 101)=? B5. What is the value of x 8 when x=27 B6. The temperature in Miami, Florida was 73° F, in Buffalo New York, the temperature was -16°. How much warmer was it in Miami than in Buffalo? B7. A byte is the amount of memory a computer uses to store one character. A byte contains 8 bits. A kilobyte is 2 10 bytes, evaluate 210 .How many bits are there in one kilobyte? Write this number in standard notation and in scientific notation. B8. The wavelength of violet light is about 0.000016 inch. A lab assistant wrote as 16 x 10-6, Is 16 x 10-6 written in scientific notation? If not, show the number should be written? (USING VARIABLES) B9. Evaluate 6m3- 5m3+8n 2 -7n2, when m=4 and n=3 B10. Write an expression using the x to represent an integer: "The sum of three consecutive integers." B11. Write an expression using the letter x to represent a multiple of five: "The sum of three consecutive multiples of five" B12. The sum of three consecutive odd integers is 87. where x is an odd integer. B13. The sum of three consecutive even integers is 102. where x is an even integer. B14. From the top of Mt. Mckinley to the floor of Death Valley is 6,250 meters. Death Valley is 86 meters below sea level. What is the height of Mt. Mckinley above sea level? B15. On a winter day in Chicago, the temperature rose 2° F every hour for 8 hours. The low temperature was -3°F. What was the temperature, t, after 8 hours? B16. Nancy's bowling handicap is 8 less than half his average. Her handicap is 62. Which of the following represents her average, a?[(62+8) x 2=a, 62÷2)÷8=a. 62÷8+2)] (INTEGERS AND REAL NUMBERS) Simplify B17. 4+(-7)+2+(-5), 144+(-317)+213, 4+17+(-29)+13+(-22)+(-3). |43-62|+|-17-3|, B18. In the St. Elias Mountains. Mt. Logan rises from 1,292 meters above sea level to 7,243 meters above sea level. How tall is Mt. Logan? B19. |-42|x|7|, (-6) x 5 x (-10) x (-4) x 0 x 2, -|-6x8|, B20. Georgia had a record low temperature of -3F°. Michigan experienced a record low that was 9 times the record low set in Georgia during the same year. What was the record low in Michigan that year? B21. (-24)÷8, (-180)÷12, |-76|÷|-4|, |-216 ÷(-6),| B22. At the end pf the year, a small firm has$2,996 in its account for bonuses. If the entire amount is equally divided among the 14 employees, how much does each one receive? B23. 4-8x2 , 96÷3÷4÷2, 3÷4x2-6÷3 [-4-8)x3]÷9, |-8-4|÷3x6+(-4), (29-17)+|-8÷3| PERCENTAGES B24. Find 3% of 80, 0.05 of 182, 83% of 166, 300% of 4, 125% of 400 B25. 48 percent of the 1,200 students at Ryan High are male. How many are female? B26. Of 220 applicants for a job, 75% were offered an initial interview, How many people were not offered an initial interview? B27 CHANGING PERCENTS TO FRACTIONS & FRACATIONS TO PERCENTS: 25%, 33 1/3, 200%, 2%, 84%, 0.0087% ⅔, 3/5, 17/20, 45/50, 1 ¼, 1/9 B28.CHANGING PERCENTS TO DECIMALS & DECIMALS TO PERCENTS: 42%, 0.3%. 8%, 175%. 34%, 0.43, 0.08, 3.0, 0.6524, 125.02 B29.(Decimal, numeral system) How to read 8760,245,295.873546 What are whole number? What are decimal fractions? B30. Change Arabic to Roman numerals: 1,5,10,50,100,1000,5000 B31.EXPONENTS: SIMPPLIFY: 42x40, 66 x 65x 64, a4b3 x a2b3, m2n3 x m3n6x m5n4,m5n3xm4n6 x m 2 n 2 B32 (SIMPLILIFY): 65 ÷ 62 x 6 4 4√ 5 x 2√ 5 9√ 6 x 4√2 √10÷√2 √25 ÷√100 , 3√12 + √3, √ 5 + 6√5 + 3√5, 6√15 + 8√15 +16√15, 6√5 +2√45, 14√2 -19√2, 10√2 - 3√8, 4√3 -2√12 33. Find the mean (averages) [18, 25, and 32] [4/9, 2/3, 5/6] [97, 102, 116, 137] 34. Find the mode(s) of each set of number. [16, 14, 12, 16, 30, and 28 ] [4, 3, 9, 2, 4, 5, and 2 ] [1, 3, 7, 4, 3, and 8] 35. A business partnership made a profit of $16,000 one year. If one partner owned four times as many shares as the other partner, how much of the profit went to the partner with the bigger share? 36. When any number is divided by 11, What is the greatest possible remainder of the resulting numbers? 37.. If x is an even number. How many values of x satisfy both of the following conditions x < 22 and x >10? 38. Chris has an average of 87 on the first 3 exams in her physics course. What grades must he get on the fourth exam so that his average on the first 4 exams is 90? A 39. If t=9 feet, What is the value of d if d=16 t2 40.. 15% of 300 = 4% of what? ___3 in__ 4 in B C 41.. 70 + (-3) 2 (-2)3 (-1)4 =? B42 5 in B43 42. Find the length of x ? _______________ 43. Point B is the center of the circle. AB=4 inches and / ABC is a right angle. The area of the shaded region is closest to [20, 25, 30, 33, 40 sq in] 44 If √x+y+A = √x+√y, find the value of A, when x = 25 and y =144 45. When 2 is added to 4 times a number, the result is 18. Which of the following equations represents this statement?[4(N+2)=18, 4N+2=18, 4(4N+2)=18,4N-2=18] 46. If the degree measures of the angles of a triangle are in the ratio 3:4:5, what is the degree measure of the smallest angle? 47. 10 percent of 90 is equivalent to ½ of what number? 48. Which of the following integers is a divisor of both 28 and 70?[4, 8, 10, 12, 14] B49.Angle ABC is bisected by line DB. If angle B50 B51 DBC=40°, what is the measure of angle ABC? 50. The area of triangle is 36. X=? 51. .The volume of the cube is 64. What is the distance from point P to point Q? 52. A stockbroker keeps 10% of the profit he makes for his client. If the customer received $8,100 from the stockbroker, what was the total profit for the transactions made by the stockbroker? [9,000 7,200 10,000 12,2500] 53. For all integers, let ® = R3 when r is an even integer. ®=R2 when R is an odd integer. What is the value of 4 - 3 ? 54.How deep will the sand be in a rectangular sand box with inside dimensions 400 centimeters long, 300 cm wide, and 20 cm deep if 480,000 cubic cm of sand is evenly distributed in it? [2, 4, 6, 8, 20 cm] 55. If the cost of 8 equally priced computer printers is $2,800, what is the total cost of 6 of the se printers? [$1,400 $1,800 $2,000 $2.100 $2,300] 56. Marbles are placed into 7 jars, one at a time. The first marble goes into jar 1, the second into jar 2, and so on until each jar has one marble. If this pattern is repeated, beginning each time with the first jar, into which jar will the 81st marble be placed? 57.A certain perfume bottle is completely filled by three perfumes, A,B,C, If there are equal amounts of A,B, and C. If there are equal amounts of A and B and four times as much C as there is A, what percent of the perfume bottle is filled by C? 58. If a watch is set at 2 p.m. and loses 1.5 minutes every 2 hours. What will lit read when it is actually 2 p.m. the next day? [1:30 1:36 1:40 1:42 1:51] 59. If Z= 4 / 3x and WZ= 4 / 6x for x ≠ 0, then W=? [2, 4, ½ ,¼, 1/8] 60. Timmy can build a dog house in 1 day. Tiffany can build the same type of house in 2 days. If they work together, how long will it take them to build 18 houses? 61. If the radius of a circle is increased by 20%, how many % increase of the area of the circle? [20%, 33%, 37.5%, 44%, can't be determined ] 62. If a=2b=5c, then the average (arithmetic mean) of a,b,and c in terms of a is [17a/30, 17a/10, 8a/3, 8a, 17a] 63. The sum of 2 numbers is 231. The larger is twice the smaller,what are the numbers? B64. The length of a rectangle is 3 times the width. If the perimeter is 80 feet, what are the dimensions? Remember to draw the figure. 65 The length of a rectangle is 6 less than twice the width. If the perimeter is 60 inches. What are the dimensions? (draw a figure) 66 A man earns $1,000 more than 3 times what his son earns. How much each earn if they both earn $29,000? 67 A piece of wire 60 inches long is cut into 2 pieces so that the larger piece is 10 inches longer than the shorter one. How long is each one? 68 Two tuna boats start from the same port at the same time, but they head in opposite directions. The faster boat travels 10 knots per hour faster than the slow boat. At the end of 8 hours, they were 272 nautical miles apart. How many miles had they each traveled? 69. The length of the larger square is twice as large as the length of the smaller square. If the sum of their perimeter is 648 cm. What is the length of each square? 70. There are 38 students in the algebra class, there are 10 less than twice as many boys as girls. How many boys and girls are there ? 71. The cost of a bottle and a cork is $1.10. If the bottle costs $1.00 more than the cork, what is the cost of each? 72. A man will left $75,000 to his wife, son and daughter in such a way that his son's share was 3 times the daughter's share, while the wife received as much as the son and daughter together. How much did each receive? (Hint: x= daughter's share) 73. (Practice Problems: Age) Mr. Ho is 25 years older than his daughter. In 10 years He'll be twice as old as his daughter. How old is each now? 74. Mr. Chu is 5 times as old as his son Alan. Five years ag0 the father was 15 times as old as his son. What are their ages now? 75.The sum of Jess's age and his pal Tim's age is 41. Six times Jesse's age 10 years ago was equal to 2 more than twice Tim's age now. How old is each now? 76. Catherine is 10 years older than her sister Hillary. In 10 years Hillary will be ¾ as old as Catherine. How old is each now? 77. Carol's age 8 years ago was 5 times Carl's age. Five years ago Carol's age was 3 times Carl's age. How old is each now? D78. Mrs. Huang is now 3 times as old as her son Jesse. Five years from now. Mrs Huang will be 7 times her son's age 5 years age. What are their age now? (Order each set of numbers from least to greatest) 79. ( 9/16, 0.575, 11/20, 12/21) (6/5, 1.25, 11/9, 7/6) (10/150,7/100, 3/50, 9/125) 80. At the family union. Reading noticed that the number of people attending could be divided into 3, 4, 5 or 6 equally groups, what are the smallest number of people? 81.Grace has $10..00. She plan to buy 4 bagels and pastries with left money, what is the maximum number of pastries she can buy? (Price list: pastries:$1.75 Bagles:$1.00) 82.Emily made bamboo fishing poles for her sister and herself. She used 3/5 of a 10-foot stick of bamboo. How long are each fishing pole? 83. There are 300 students in the eighth grade. The ratio of girls to boys is 3 to 2. What percent of the students are boys? 84. Richard, Roger, and Jesse collected cans for a good drive. The sum of Richard and Roger's are less than 50. The sum of Roger and Jesse's are more than 50. Which statement is right?[Jesse collected more than Roger, Roger collected more than Jesse, Jesse more than Richard, Roger and Jesse are the same] 85.A car travel 187 miles in 3 hours and 24 minutes, what was the average speed in miles per hour? 86. Teresa has 3 jars of marbles that contain x marbles, and 2 jars that contain y marbles, what expression represent the total number that she has? 87.In a group of 40 people, 32 are over or equal 5 feet 7 inches in height, what is the probability that the person's height under 5 feet 7 inches? 88. In the first race of the season. Jason swam his event in 5 minutes and 15 seconds. His goal was 2/3 of the time. How long will it take for him at the end of the season? B89 How much fencing is needed to enclose the garden? B90.What is the relationship between angles A and C? [complementary, supplementary, congruent or vertical angles] B91. Jefferey surveyed his classmates to see how many of them owned bicycles, in-line skates, and scooter. He made this Venn Diagram to display the results. What is the ratio of classmates who own all three? [5/24, 5/37, 5/42, 5/47] Jefferey's survey B89 Scale:1 centimeter=1 meter bicycle in line skate 9 7 scooter B92. For all numbers, (a-b)(b-c)-(b-a)(c-b)=? _________ __________ 93. If 1 / 2N + 1 / 2N = ¼, then N=? P S U W Q __________ __________ 94. If x/z = k and y/z = k-1, then x=? 95. The map showing the station s and connecting subway line for a city's subway system. If a person wishes to travel by subway from station P to Q, without passing through any station more then once, which of the following must be true?[If he passes through T, he must next pass through S, If he passes through S, he must next pass through U, If he passes through R, he can't later pass through T, If he passes through V, he can't later pass through W] 96. In a certain game, a person's age is multiplied by 2 and then the product is divided by 3, If the result of performing the operations on John's age is 12, what is John's age? 97. A helper must lad 38 bricks onto a truck., Given that she can carry at most 4 bricks at a time, what is the fewest number of trips she must make to move all of the bricks from the brick pile onto the truck? 98. N is a positive integer. If N is a multiple of 6 and a multiple of 9, what is the least possible value of N? 99. For any positive integer k, = k2 - k. What is the value of ___ ? 100. A square piece of land that is divided into 9 smaller square lots. The shaded portion is a railroad right-of-way. If the area of the shaded portion is 5 square miles. What is the area, in square miles, of the entire piece of land? [9,10,13,18,36] (Quantitative Comparisons) Mark A if the quantity in Column A is greater Mark B if the quantity in Column B is greater Mark C if the two quantity are equal Mark D if the relationship can't be determined from the information given Note: Information that applies to both quantities is centered above both columns A symbol that appears in both columns means the same thing in both columns Column A Column B______________________ 10% of x =15 (A) 20 10% of 2x (B) x 2 + y 2 (x+y)2 A E B E is any point on side AB (C) Area of shaded portion Area of unshaded portion of square ABCD of square ABCD

OpenStudy (anonymous):

?? I'm not really sure that helps me with my question, haha. Sorry.